# On the KG-constrained general disformal transformation of the Einstein-Hilbert action

### v2

@@ Abstract
Although with great successes in explaining phenomena and natural behaviour involving the Universe or a part thereof, the General Theory of Relativity is far from a complete theory. There are plenty of avenues to investigate its extensions or generalisations that, in the future, could pave the way for a much deeper understanding and possibly solutions of current problems involving the Universe; e.g., dark matter, dark energy, fine-tuning, etc. Focusing on its extension within the framework of scalar tensor theory, we investigate the transformation of the Einstein-Hilbert action under the general disformal transformation coupled with a set of triple KG constraints. The idea is that the general disformal transformation leads to an extremely complicated action necessitating a set of constraints to tame the resulting form. Motivated by previous studies on the invertibility and invariance of the massless Klein-Gordon equation under the general disformal transformation, we identify three constraints that significantly simplifies the transformed Einstein-Hilbert action. In four spacetime dimensions, we find that the transformed action is a sum of the original action and a disformal contribution involving a six-term Lagrangian that includes the Ricci tensor coupled to a sum of derivatives of scalar fields and kinetic terms. In three spacetime dimensions, the disformal contribution becomes a five-term Lagrangian. Lastly, in two spacetime dimensions, the Einstein-Hilbert action is invariant under the constrained disformal transformation.
@@ 1 Introduction
Disformal transformation is a metric transformation generalising the more familiar conformal transformation in general relativity. Bekenstein introduced disformal transformation in the early 1990s to relate the physical metric describing Riemannian geometry, to the gravitational metric, within the context of Finslerian geometry.[1] Since then, disformal transformation found numerous applications beyond the initial framework within which the transformation was introduced. Within two decades after the publication of Bekenstein’s work, there were several studies on symmetries of primordial cosmological perturbations [^2] For studies on the conformal transformation involving primordial cosmological perturbation, interested readers may refer to Refs. [6, 7, 8, 9, 10]. [2, 3, 4, 5] under disformal transformation, disformal and Higgs cosmic inflations [11, 12, 13, 14, 15], invariance of field equations of motion [16, 17, 18, 19], variations of actions in scalar tensor theories [20, 21, 22, 23, 24], etc. More recently, researches on disformal transformation involve theoretical investigations of wormholes [25], hairy black holes [26], mimetic gravity [27], and “most” general action forms subject to invertibility in relation to Ostrogradsky instability [28, 29, 30], to name a few.
@@
Having said this, we see more than one way to impose conditions on the functionals in the general disformal transformation above if the aim is a relatively ‘simple’ resulting action; that is, one with small number of terms and possible instabilities. Knowing the time and space limitations in this paper, we find it prudent to use one good set of constraints and leave other possible sets of constraints for future studies. In this study, we look no further and employ the constraints in our earlier work in Ref. [17] involving the massless Klein-Gordon equation [^6] This work was inspired by Ref. [16] where the massless Klein-Gordon equation was found to be invariant under the (less general) Bekenstein’s disformal transformation under some constraint on the disformal factor.. Based on this study, the massless Klein-Gordon equation for scalar fields and correspondingly, the Klein-Gordon action, remain invariant under the general disformal transformation given by (2) if the disformal factors $C$ and $D$ take on a special form in relation to the conformal factor $A$; see Sec. 2 below.
The idea for this choice of constraint in the current work is at least two-fold. First, the Klein-Gordon action and the Einstein-Hilbert action both belong as (two distinct) special cases of the Horndeski action [37]. As such, it is interesting to know if the general disformal transformation of the Einstein-Hilbert action, subject to the same constraints, is invariant or at least reduces to a subset of the Horndeski action, similar to that in Ref. [21]. In hindsight, even if the second scenario is not completely attainable due to the presence of a few higher order derivative terms, we would find it odd to have two ‘subsets’ of the Horndeski action to have different constrained (symmetry) transformations [^7] We leave it for future work to investigate such a possible unusual scenario.. Second, the constrained disformal transformation in Ref. [17] that yields invariant massless Klein-Gordon equation is invertible, in general. Consequently, using the same constrained transformation frees us from the problem of invertibility (of the transformation itself) that could lead to additional possible instabilities and issues involving ghost terms and the primordial cosmological perturbations, within the much larger framework of Horndeski theory and extensions thereof.
@@ 2 Invariance of the massless Klein-Gordon equation under disformal transformation
To proceed, one may start with the expansion of the double hatted covariant derivatives in the expression for ^$\widehat{\square}\phi$ as ^$\widehat{\nabla}_{\mu}\widehat{\nabla}_{\nu}\phi=\phi_{;\nu,\mu}-\widehat{\Gamma}^{\alpha}_{\mu\nu}\phi_{;\alpha},$ where the hatted Christoffel symbol, $\widehat{\Gamma}^{\alpha}_{\mu\nu},$ simply follows the form of the original one, $\Gamma^{\alpha}_{\mu\nu}$; that is,
@@
$\Phi_{(\mu;\nu)}$ $\equiv\tfrac{1}{2}(\Phi_{\mu;\nu}+\Phi_{\nu;\mu}),$
$\Phi_{[\mu;\nu]}$ $\equiv\tfrac{1}{2}(\Phi_{\mu;\nu}-\Phi_{\nu;\mu}).$ (8)
@@ 3 KG-Constrained disformal transformation of the Ricci tensor
Note that the last two lines above contain third-order derivative terms involving $\Phi_{\mu}$ and ultimately, the field variables $(\phi,X,Y,Z)$. In fact, $F^{\alpha}_{\mu\nu;\alpha}$ is the only source of these higher order derivative terms in the calculation for $\widehat{R}_{\mu\nu}$, coming from the covariant derivative of $F^{\alpha}_{\mu\nu}$. That is, given the set of KG constraints without which the term $F^{\alpha}_{\mu\alpha;\mu}$ could have added more third order derivative terms. With this being said, at first glance, these terms may look quite problematic, at least if we consider working on the field variables $(\phi,X,Y,Z)$. However, because the third-order derivative factors are multiplied with at most first-order derivative factors,131313e.g., $\Phi_{[\beta;\nu];\alpha}$ is multiplied with $\Phi_{\mu}$ and functional of $A$. then we may simply perform integration by parts and effectively transfer one of the derivative indices in the third-order derivative terms to its ‘neighbours’. We deal with these terms in the next section upon derivation of the disformal Ricci scalar.
@@ 5 The KG-constrained disformally transformed Einstein-Hilbert action in $n=4,3,2$ dimensions
With this being said, one may wonder if in the ‘forest’ of terms in the full expansion of $S^{(4)}$ we may find fragments of the Horndeski action [37]:
@@
where $G_{2}$ to $G_{5}$ are all functionals of $(\phi,X)$ and $G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R$ is the Einstein tensor. It may be recalled that the special disformal transformation of the Einstein-Hilbert action yields a special case of the Horndeski action covering the sub-Lagrangians $\mathcal{L}_{2},\mathcal{L}_{3},$ and $\mathcal{L}_{4}$ [21], consistent with the hierarchical propagation of terms observed in Ref. [20]. For our case here, we can also identify terms belonging to the Horndeski sub-Lagrangians but only for $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ with all other numerous terms in the action being beyond-Horndeski terms.
@@
This implies that the Einstein-Hilbert action is in general, not invariant under conformal transformation. Nevertheless, for $n=2$,
@@ 6 Summary and concluding remarks
As it turns out, the general disformal transformation of the $n$-dimensional Einstein-Hilbert action, subject to the triple KG constraints, leads to a sum of the original action and a disformal contribution involving only a six-term Lagrangian that includes the Ricci tensor coupled to a sum of derivatives of scalar fields and kinetic terms. Although the transformation does not lead to invariance in general spacetime dimensions, the number of ‘extraneous’ terms at the level of $\Phi_{\mu}$ is significantly reduced when compared to that for unconstrained general disformal transformation. Furthermore, we identify two of the four Horndeski sub-Lagrangians resulting from the transformation. In four spacetime dimensions, we find the number of disformal contribution terms in the Lagrangian to be the same as that for $n>5$. This decreases to a five-term Lagrangian from the original six-term disformal contributions, in three dimensions. Interestingly, in two spacetime dimensions, the Einstein-Hilbert action is invariant under the constrained disformal transformation!

### v3

@@ Abstract
Although with great successes in explaining phenomena and natural behaviour involving the Universe or a part thereof, the General Theory of Relativity is far from a complete theory. Focusing on its extension within the framework of scalar tensor theory, we investigate the transformation of the Einstein-Hilbert action under the general disformal transformation coupled with a set of triple KG constraints. The idea is that the general disformal transformation leads to an extremely complicated action necessitating a set of constraints to tame the resulting form. Motivated by previous studies on the invertibility and invariance of the massless Klein-Gordon equation under the general disformal transformation, we identify three constraints that significantly simplifies the transformed Einstein-Hilbert action. In four spacetime dimensions, we find that the transformed action is a sum of the original action and a disformal contribution involving a six-term Lagrangian that includes the Ricci tensor coupled to a sum of derivatives of scalar fields and kinetic terms. In three spacetime dimensions, the disformal contribution becomes a five-term Lagrangian. Lastly, in two spacetime dimensions, the Einstein-Hilbert action is invariant under the constrained disformal transformation.
@@ 1 Introduction
Disformal transformation is a metric transformation generalising the more familiar conformal transformation in general relativity. Whereas the conformal transformation involves scaling of the metric, $g_{\mu\nu}$, by a (scalar) field-dependent conformal factor, $A(\phi)$, that is, $g_{\mu\nu}\rightarrow\widehat{g}_{\mu\nu}=A(\phi)g_{\mu\nu}$, the original disformal transformation due to Bekenstein [1], includes a direction-dependent variation of the metric in the transformation. In particular, $\widehat{g}_{\mu\nu}=A(\phi,X)g_{\mu\nu}+B(\phi,X)\phi_{;\mu}\phi_{;\nu}$, where $B$ is the disformal factor and the semicolon ‘;’ denotes covariant derivative; i.e., $\phi_{;\mu}=\nabla_{\mu}\phi$. The disformal part, $B(\phi,X)\phi_{;\mu}\phi_{;\nu}$, serves to deform the metric along the direction of variation of the scalar field. In addition to this, the generalisation includes dependence on the kinetic term, $X\equiv-(1/2)g^{;\mu\nu}\phi_{;\mu}\phi_{;\nu}$, which also happens to encode the (inverse) metric. In perspective then, the Bekenstein’s disformal transformation is a two-fold generalisation of the usual conformal transformation.
Because of this, it is not straightforward to have a relatively simple physical intuition about the disformal transformation beyond our concise description above, in contrast to that for the conformal transformation. Nonetheless, we can take a few steps forward in its physical interpretation [^2] Readers interested in a more thorough discussion of the physical interpretation of, constraints on, and value of the disformal transformation in relation to the conformal transformation and the more general Scalar-tensor theories, may refer to Ref. [2] and the references therein. For numerous applications of disformal transformation in relation to Horndeski action, see Ref. [3]. by switching off the dependence on $X$ above to yield the special disformal transformation; i.e., $\widehat{g}_{\mu\nu}=A(\phi)g_{\mu\nu}+B(\phi)\phi_{;\mu}\phi_{;\nu}$. When $B=0$, the conformal transformation scales the metric but preserves the angles between tangent vectors on a manifold that the metric describes. Physically, this means that the light cones associated with the manifold remain the same modulo scaling and the underlying causal structure is preserved. When $B\neq 0$, the disformal transformation can widen or narrow the light cones. Certainly, the latter possibility can lead to undesirable acausal structures. However, with suitable constraint(s) in the conformal and disformal factors (see Ref. [2]), disformal transformation can give rise to a richer set of structures where the light cones may ‘fold’ within the bounds of causality.
Beyond this, when the conformal and disformal factors are no longer dependent on $\phi$ alone, and additional kinetic-term dependencies may come into play, we are welcomed with far richer set of possible structures and flexibility that are not found in conformal transformation alone. Bekenstein, who introduced disformal transformation, exploited this greater ‘flexibility’ in the 1990s to relate the physical metric describing Riemannian geometry, to the gravitational metric, within the context of Finslerian geometry [1]. Since then, unbounded by the limitations of conformal transformation alone, disformal transformation found numerous applications beyond the initial framework within which the transformation was introduced. Within two decades after the publication of Bekenstein’s work, there were several studies on symmetries of primordial cosmological perturbations [^3] For studies on the conformal transformation involving primordial cosmological perturbation, interested readers may refer to Refs. [8, 9, 10, 11, 12]. [4, 5, 6, 7] under disformal transformation, disformal and Higgs cosmic inflations [13, 14, 15, 16, 17], invariance of field equations of motion [18, 19, 20, 21], variations of actions in scalar tensor theories [2, 3, 22, 23, 24], etc. More recently, researches on disformal transformation involve theoretical investigations of wormholes [25], hairy black holes [26], mimetic gravity [27], nonminimally coupled Proca theory [28], and “most” general action forms subject to invertibility in relation to Ostrogradsky instability [29, 30, 31], to name a few.
@@
Having said this, we see more than one way to impose conditions on the functionals in the general disformal transformation above if the aim is a relatively ‘simple’ resulting action; that is, one with small number of terms and possible instabilities. Knowing the time and space limitations in this paper, we find it prudent to use one good set of constraints and leave other possible sets of constraints for future studies. In this study, we look no further and employ the constraints in our earlier work in Ref. [19] involving the massless Klein-Gordon equation [^7] This work was inspired by Ref. [18] where the massless Klein-Gordon equation was found to be invariant under the (less general) Bekenstein’s disformal transformation under some constraint on the disformal factor.. Based on this study, the massless Klein-Gordon equation for scalar fields and correspondingly, the Klein-Gordon action, remain invariant under the general disformal transformation given by (2) if the disformal factors $C$ and $D$ take on a special form in relation to the conformal factor $A$; see (16), (17), and (18) in Sec. 2 below.
The idea for this choice of set of constraints in the current work is at least two-fold. First, the Klein-Gordon action and the Einstein-Hilbert action both belong as (two distinct) special cases of the Horndeski action [38]. As such, it is interesting to know if the general disformal transformation of the Einstein-Hilbert action, subject to the same constraints, is invariant or at least reduces to a subset of the Horndeski action, similar to that in Ref. [22]. In hindsight, even if the second scenario is not completely attainable due to the presence of a few higher order derivative terms, we would find it odd to have two ‘subsets’ of the Horndeski action to have different constrained (symmetry) transformations [^8] We leave it for future work to investigate such a possible unusual scenario.. Second, the constrained disformal transformation in Ref. [19] that yields invariant massless Klein-Gordon equation is invertible, in general [^9] Interested readers may refer to Ref. [30] for a more general treatment of invertibility.. Consequently, using the same constrained transformation frees us from the problem of invertibility (of the transformation itself) that could lead to additional possible instabilities and issues involving ghost terms and the primordial cosmological perturbations, within the much larger framework of Horndeski theory and extensions thereof.
@@ 2 Invariance of the massless Klein-Gordon equation under disformal transformation
To proceed, one may start with the expansion of the hatted double covariant derivatives in the expression for ^$\widehat{\square}\phi$ as ^$\widehat{\nabla}_{\mu}\widehat{\nabla}_{\nu}\phi=\phi_{;\nu,\mu}-\widehat{\Gamma}^{\alpha}_{\mu\nu}\phi_{;\alpha},$ where the hatted Christoffel symbol, $\widehat{\Gamma}^{\alpha}_{\mu\nu},$ simply follows the form of the original one, $\Gamma^{\alpha}_{\mu\nu}$; that is,
@@
$\Phi_{(\mu;\nu)}$ $\equiv\tfrac{1}{2}(\Phi_{\mu;\nu}+\Phi_{\nu;\mu})\quad$
$\text{and}\quad\Phi_{[\mu;\nu]}$ $\equiv\tfrac{1}{2}(\Phi_{\mu;\nu}-\Phi_{\nu;\mu}).$ (8)
@@ 3 KG-Constrained disformal transformation of the Ricci tensor
Note that the last two lines above contain third-order derivative terms involving $\Phi_{\mu}$ and ultimately, the field variables $(\phi,X,Y,Z)$. In fact, $F^{\alpha}_{\mu\nu;\alpha}$ is the only source of these higher order derivative terms in the calculation for $\widehat{R}_{\mu\nu}$, coming from the covariant derivative of $F^{\alpha}_{\mu\nu}$. That is, given the set of KG constraints without which the term $F^{\alpha}_{\mu\alpha;\mu}$ could have added more third order derivative terms. With this being said, at first glance, these terms may look quite problematic, at least if we consider working on the field variables $(\phi,X,Y,Z)$. However, because the third-order derivative factors are multiplied with at most first-order derivative factors,151515e.g., $\Phi_{[\beta;\nu];\alpha}$ is multiplied with $\Phi_{\mu}$ and functional of $A$. then we may simply perform integration by parts and effectively transfer one of the derivative indices in the third-order derivative terms to its ‘neighbours’. We deal with these terms in the next section upon derivation of the disformal Ricci scalar.
@@ 4 Klein-Gordon constrained disformal transformation of the Einstein-Hilbert action
Before we leave this section and explore in the next section the special cases of the resulting action, it is worth reflecting at this point about the physical degrees of freedom involved in the current work, in relation to a related studiy mentioned above involving the disformal transformation of the action. In Ref. [2], the authors started with the Horndeski action itself and perform special disformal transformation. Knowing that this transformation is invertible and that the resulting form is the same as the original one, our insight is that there is effectively no new physical degrees of freedom ‘generated’. Within the framework of inflationary cosmology for instance, such transformation leaves the scalar and tensor primordial cosmological perturbations invariant at least in the superhorizon limit; see for instance, Refs. [4, 5, 6, 7, 37].
In the current work however, the original action is simply a special case of the Horndeski action, namely, the Einstein-Hilbert action. In the absence of additional terms in the original action that could possibly lead to some cancellation and owing to the generality and complexity of the disformal transformation, we gain a more general action that includes two sub-Lagrangians of the Horndeski action and several beyond-Horndeski terms. These new Horndeski and beyond-Horndeski terms are suggestive of new physical degrees of freedom rooted in the conformal and disformal factors. However, we have the insight that as the disformal factors are linked to the ‘free’ conformal factor by the set of KG constraints, that is, both disformal factors are expressible in terms of the conformal factor, we may only have at most one additional physical degree of freedom. This is in addition to the fact that from the perspective of single scalar field Scalar-Tensor theory, we have ultimately introduced only one new field, namely, $\phi$. Beyond this insight, as to the nature and full accounting of this possible new physical degree of freedom, we see further needs for more thorough analyses and in-depth calculations. We hope to address this matter in our future studies.
@@ 5 The KG-constrained disformally transformed Einstein-Hilbert action in $n=4,3,2$ dimensions
With this being said, one may wonder if in the ‘forest’ of terms in the full expansion of $S^{(4)}$ we may find fragments of the Horndeski action [38]—the most general gravitational action within the framework of Scalar Tensor theory in four dimensions involving the metric tensor and a scalar field, whose equations of motion are at most second order. It reads
@@
where $G_{2}$ to $G_{5}$ are all functionals of $(\phi,X)$ and $G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R$ is the Einstein tensor [^17] It encodes many gravitational theories including General Relativity and Brans-Dicke theory, to name just a couple of them. Indeed, when $G_{4}\sim\text{const.}$. and all other $G_{i}$’s are vanishing, we get the (geometric part) of the Einstein-Hilbert action. On the other hand, if $G_{4}\sim\phi$ and $G_{2}\sim X/\phi$ with $G_{3},G_{5}=0$, we get the Brans-Dicke theory[53].. It may be recalled that the special disformal transformation of the Einstein-Hilbert action yields a special case of the Horndeski action covering the sub-Lagrangians $\mathcal{L}_{2},\mathcal{L}_{3},$ and $\mathcal{L}_{4}$ [22], consistent with the hierarchical propagation of terms observed in Ref. [2]. For our case here, we can also identify terms belonging to the Horndeski sub-Lagrangians but only for $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ with all other numerous terms in the action being beyond-Horndeski terms.
@@
This implies that the Einstein-Hilbert action is in general, not invariant under conformal transformation. Nevertheless [^19] On a related note, the fact that the Einstein-Hilbert action is topological in nature in $n=2$ dimensions may lead one to introduce a (dilaton) field in a form coupled to $R$ and a corresponding (linear) potential as in the case of Jackiw-Teitelboim gravity[54, 55]. This addresses the problem of the suppression of fluctuation in quantum gravity for $n=2$ dimensions., for $n=2$,
@@ 6 Summary and concluding remarks
As it turns out, the general disformal transformation of the $n$-dimensional Einstein-Hilbert action, subject to the triple KG constraints, leads to a sum of the original action and a disformal contribution involving only a six-term Lagrangian that includes the Ricci tensor coupled to a sum of derivatives of scalar fields and kinetic terms. Although the transformation does not lead to invariance in general spacetime dimensions, the number of ‘extraneous’ terms at the level of $\Phi_{\mu}$ is significantly reduced when compared to that for unconstrained general disformal transformation. Furthermore, we identify two of the four Horndeski sub-Lagrangians resulting from the transformation. In four spacetime dimensions, we find the number of disformal contribution terms in the Lagrangian to be the same as that for $n>4$. This decreases to a five-term Lagrangian from the original six-term disformal contributions, in three dimensions. Interestingly, in two spacetime dimensions, the Einstein-Hilbert action is invariant under the constrained disformal transformation!